# Investigating the Robustness of LLMs on Math Word Problems

### v2

@@ Abstract
Large Language Models (LLMs) excel at various tasks, including solving math word problems (MWPs), but struggle with real-world problems containing irrelevant information. To address this, we propose a prompting framework that generates adversarial variants of MWPs by adding irrelevant variables. We introduce a dataset, ProbleMathic, containing both adversarial and non-adversarial MWPs. Our experiments reveal that LLMs are susceptible to distraction by numerical noise, resulting in an average relative performance drop of on adversarial MWPs. To mitigate this, we fine-tune LLMs (Llama-2, Mistral) on the adversarial samples from our dataset. Fine-tuning on adversarial training instances improves performance on adversarial MWPs by , indicating increased robustness to noise and better ability to identify relevant data for reasoning. Finally, to assess the generalizability of our prompting framework, we introduce GSM-8K-Adv, an adversarial variant of the GSM-8K benchmark. LLMs continue to struggle when faced with adversarial information, reducing performance by up to .
@@ 1 Introduction
Additionally, we compare our approach with others in the domain of adversarial data generation in §B.
@@ 2.1 Adversarial data generation
We propose a constrained additive approach, wherein we add numerical information (Structural Invariance) to the original MWP. We use our hand-crafted set of MWPs as seed samples and augment them in accordance with the following constraints: (1) The added variables must not be related to or derived from the existing variables in the passage. (2) The added variables must not share the same physical unit as any of the original variables. (3) The augmented text should not add any numerical information about existing variables that did not exist in the original passage. This gives us a set of adversarial MWPs containing variables that do not feature in their solution equations. Hence, this allows us to evaluate the model’s ability to identify relevant information and demonstrate reasoning ability beyond logically connecting all numerical entities in the input in order to produce a solution. The prompting framework used to adversarially augment the seed samples can be found in §E. We also compare our adversarial data generation approach with other contemporaneous approaches in §B.
@@ 2.2 Dataset
ProbleMathic is split into Simple and Complex problem sets, containing both adversarial and non-adversarial samples. The Simple set consists of MWPs requiring a single mathematical operation to solve. Number of samples in train and test splits are 1313 and 60 respectively. These samples are aggregated from existing sources detailed in §C. The Complex set contains MWPs involving more than one mathematical operation. Number of samples in train and test splits are 613 and 73 respectively. Table 1 highlights some examples of train and test set of the dataset. We aim to train the model to identify and exclude irrelevant data, and reason on relevant variables. To this end, we use GPT-4 Achiam et al. (2023) to generate solution explanations for all training samples. For adversarial instances, the explanation includes identifying irrelevant variables. The adversarial test instances are hand-crafted to eschew any data leakage (§C.1).
@@ 3.2.2 Fine-tuning
Table 4 presents LLM fine-tuning results across various training data scenarios: Og (non-adversarial), adversarial, and a combined set. Performance is evaluated on Og and adversarial samples from both Simple and Complex problem sets. Llama-2 (13B) is the best performer, likely due to its large size. The optimal fine-tuning setting, highlighted in varies across test sets. While similar samples yield the best performance on Og Simple problems, training on adversarial samples boosts average performance on adversarial Simple problems and both Og and adversarial Complex problems by 8.19%, 7.07%, and 1.78%, respectively. Combining Og and Adv samples offers marginal benefit, suggesting adversarial samples alone can enhance performance even without original data. We also present fine-tuning results with these models on the MultiArith Roy and Roth (2016) dataset in §A.
@@ 5 Conclusion
Appendix A Results on MultiArith dataset
@@
Appendix B Comparison with other adversarial generation frameworks
@@
B.1 Comparison with "Adversarial Math Word Problem Generation"
@@
Appendix C Curating ProbleMathic
@@
C.1 Human-crafted test samples
@@
Appendix D Human quality evaluation
@@
Appendix E Prompting Framework for Adversarial Data Generation

### v3

# Cutting Through the Noise: Boosting LLM Performance on Math Word Problems

@@ Abstract
Large Language Models (LLMs) excel at various tasks, including solving math word problems (MWPs), but struggle with real-world problems containing irrelevant information.
To address this, we propose a prompting framework that generates adversarial variants of MWPs by adding irrelevant variables.
We introduce a dataset, ProbleMathic, containing both adversarial and non-adversarial MWPs.
Our experiments reveal that LLMs are susceptible to distraction by numerical noise, resulting in an average relative performance drop of on adversarial MWPs.
To mitigate this, we fine-tune LLMs (Llama-2, Mistral) on the adversarial samples from our dataset.
Fine-tuning on adversarial training instances improves performance on adversarial MWPs by , indicating increased robustness to noise and improved ability to identify relevant data for reasoning.
Finally, to assess the generalizability of our prompting framework, we introduce GSM-8K-Adv, an adversarial variant of the GSM-8K benchmark.
LLMs continue to struggle when faced with adversarial information, reducing performance by up to .
111 Currently in Microsoft Equal Contribution
Currently in Amazon (The work was done prior to joining Amazon) Currently in Google DeepMind
Codebase: https://github.com/him1411/problemathic
Dataset: https://huggingface.co/datasets/him1411/problemathic
@@ 1 Introduction
Additionally, we compare our approach with others in the domain of adversarial data generation in §C.
@@ 2.1 Adversarial data generation
We propose a constrained additive approach, wherein we add numerical information (Structural Invariance) to the original MWP. We use our hand-crafted set of MWPs as seed samples and augment them in accordance with the following constraints: (1) The added variables must not be related to or derived from the existing variables in the passage. (2) The added variables must not share the same physical unit as any of the original variables. (3) The augmented text should not add any numerical information about existing variables that did not exist in the original passage. This gives us a set of adversarial MWPs containing variables that do not feature in their solution equations. Hence, this allows us to evaluate the model’s ability to identify relevant information and demonstrate reasoning ability beyond logically connecting all numerical entities in the input in order to produce a solution. The prompting framework used to adversarially augment the seed samples can be found in §F. We also compare our adversarial data generation with other contemporaneous approaches (§C).
@@ 2.2 Dataset
ProbleMathic is split into Simple and Complex problem sets, containing both adversarial and non-adversarial samples. The Simple set consists of MWPs requiring a single mathematical operation to solve. Number of samples in train and test splits are 1313 and 60 respectively. These samples are aggregated from existing sources detailed in §D. The Complex set contains MWPs involving more than one mathematical operation. Number of samples in train and test splits are 613 and 73 respectively. Table 1 highlights some examples of train and test set of the dataset. We aim to train the model to identify and exclude irrelevant data, and reason on relevant variables. To this end, we use GPT-4 Achiam et al. (2023) to generate solution explanations for all training samples. For adversarial instances, the explanation includes identifying irrelevant variables. The adversarial test instances are hand-crafted to eschew any data leakage (§D.1).
@@ 3.2.2 Fine-tuning
Table 4 presents LLM fine-tuning results across various training data scenarios: Og (non-adversarial), adversarial, and a combined set. Performance is evaluated on Og and adversarial samples from both Simple and Complex problem sets. Llama-2 (13B) is the best performer, likely due to its large size. The optimal fine-tuning setting, highlighted in varies across test sets. While similar samples yield the best performance on Og Simple problems, training on adversarial samples boosts average performance on adversarial Simple problems and both Og and adversarial Complex problems by 8.19%, 7.07%, and 1.78%, respectively. Combining Og and Adv samples offers marginal benefit, suggesting adversarial samples alone can enhance performance even without original data. We also present fine-tuning results with these models on the MultiArith Roy and Roth (2016) dataset in §B.
@@ 5 Conclusion
Appendix A Hyperparameters
GPU: 8xNvidia Tesla V100, Train Batch Size: 4. Gradient Accumulation Steps: 8, Initial learning rate: 5e-5, Num of Epochs: 2
Appendix B Results on MultiArith dataset
@@
Appendix C Comparison with other adversarial generation frameworks
@@
C.1 Comparison with "Adversarial Math Word Problem Generation"
@@
Appendix D Curating ProbleMathic
@@
D.1 Human-crafted test samples
@@
Appendix E Human quality evaluation
@@
Appendix F Prompting Framework for Adversarial Data Generation